# Can Uncertainty Quantification Enable Better Learning-based Index Tuning?

### v1

@@ 1 Introduction
(1) Poor stability: Learning-based models typically outperform what-if tools on data similar to their training sets, but they can generate highly erroneous results when query patterns shift or training data is insufficient. As shown in Figure 1, we compare a learning-based BE method with a what-if tool on varying proportions of out-of-domain (OOD) samples. When the proportion reaches around 25%, the mean errors of two methods are comparable, but the mean error of the learning-based method becomes larger than the what-if tool as the OOD sample proportion increases.
@@
Due to the poor stability of learning-based models, some estimation results will inevitably be highly inaccurate and unreliable, leading the index advisor to select suboptimal indexes. Furthermore, the lack of interpretability prevents us from determining when or why the model produces unreliable results, making it difficult to fundamentally address the stability issue. In this paper, we propose a different approach by quantifying the uncertainty of the model’s results through specific metrics. When the model’s estimate for a test sample shows high uncertainty, the result is more likely to significantly deviate from the true value. This technique, known as uncertainty quantification (UQ), has been widely studied in fields where model errors have serious consequences, such as medical image detection [8], autonomous driving [9], and drug analysis [10].
@@
(1) We conducted a systematic study of the sources of uncertainty in learning-based BE models within the context of index tuning. Based on the characteristics of these tasks, we proposed a novel hybrid UQ method that allows for a flexible trade-off between efficiency and accuracy, depending on the specific characteristics of the BE task (Section 4). Our experiments demonstrate that the proposed UQ method outperforms general-purpose UQ methods in five out of six scenarios.
(2) We introduced Beauty, the first uncertainty-aware BE framework, which combines the strengths of learning-based models with what-if tools, enabling uncertainty-driven model updates. Our experiments across six benchmarks with varying budgets show that the Beauty framework significantly improves index tuning performance (Section 6).
(3) In our experiments, we applied existing UQ methods to various learning-based BE models, creating a total of 16 different models (Section 5). We thoroughly evaluated the impact of these methods on model accuracy and efficiency, confirming the superiority of our proposed UQ method.
2 Background
2.1 Index Advisors
@@
2.2 Learning-based Benefit Estimation (BE) Models
The benefit estimator plays a crucial role in index tuning as its efficiency directly impacts the overall performance of index tuning, and its accuracy affects the behavior of the index configuration enumerator [15, 20, 21]. Traditionally, what-if tools have been used to create hypothetical indexes, and the query optimizer is employed to estimate the total cost $c(W,\mathcal{I})$. However, due to their inefficiency and inaccuracy, particularly, for complex queries, machine learning (ML) techniques have been recently introduced to develop faster and more accurate BE models. One approach such as DISTILL [1] involves training multiple lightweight BE models for each query template to collectively estimate benefits. The other approach is to train a single complex BE model using historical query execution data, either to completely replace the what-if tool or to adjust the outputs of the what-if tool. The latter approach is the main focus of our study. LIB [7] and AMA [6] are notable examples of this approach. As shown in Figure 3, the typical architecture of a learning-based BE model consists of three modules.
@@ 2.2 Learning-based Benefit Estimation (BE) Models
2.3 Uncertainty in Machine Learning (ML) Models
Uncertainty is the quality or state of being uncertain, which arises in every stage of a ML task, including data acquisition, feature encoding, model design, model training, and model inference [22]. Naturally, uncertainty is inherent in ML-based index tuning, which will be explained in subsection 4.2. As illustrated in Figure 4, there are two different types of uncertainty, often referred to as aleatoric uncertainty and epistemic uncertainty [23].
@@ 2.3 Uncertainty in Machine Learning (ML) Models
Uncertainty Quantification (UQ): A number of model-independent and model-dependent methods have been proposed to quantify the two types of uncertainty. The first category includes Bayesian neural networks (BNNs) [25, 26], Monte Carlo dropout (MCD) methods [27], and ensemble methods [28, 29], which do not rely on the specific models in which the uncertainty is quantified. The second category incorporates UQ into model-specific structures. For example, HSTO [30] replaces the multi-head attention mechanism with hierarchical stochastic self-attention to achieve UQ. In this paper, we will reproduce these methods in the experimental evaluation and compare them with our proposed hybrid UQ approach.
3 Problem
Since collecting training data is costly, ML models inevitably encounter out-of-distribution (OOD) data, leading to unreliable outputs. The unreliable results produced by learning-based benefit estimators can lead index advisors to select inappropriate indexes, significantly degrading database performance. Therefore, as discussed in section 1, we will explore how UQ techniques can be leveraged to determine whether the results produced by learning-based BE models are reliable and how to integrate uncertainty-aware benefit estimators into index advisors. Specifically, this paper addresses the following three key research questions:
Q1: How can the uncertainty in a learning-based BE model be precisely quantified? Accurately quantifying the uncertainty in a learning-based BE model presents a significant challenge, and directly applying the existing methods often yields suboptimal performance. As shown in Figure 5, on a OOD dataset, model error and model uncertainty are generally not highly linearly correlated, that is, a large model error may not mean large model uncertainty. In this paper, the first problem is how to precisely quantify the uncertainty in a learning-based BE model. Rather than applying the existing general-purpose UQ methods directly, we propose a hybrid quantification method in section 4 tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods.
@@ 3 Problem
Q2: What impacts does UQ have on a BE model? In section 5, we will conduct an empirical study to understand the impacts of UQ on BE models by evaluating the accuracy, training cost, and inference cost of the BE models enhanced with UQ. In subsection 4.4, we will also explore a new method for determining when a BE model needs to be updated based on quantified uncertainty.
Q3: What impacts does a UQ-enhanced BE model have on the index advisor? In section 6, we will evaluate 8 UQ-enhanced BE models with 6 workloads across 36 scenarios to investigate the effects of the accuracy and efficiency of UQ on the index advisor.
Due to space constraints, we focus exclusively on BE models based on neural networks (NNs). NN-based models have been receiving increasing attention and generally achieve higher accuracy. The approach proposed in this paper can be easily extended to non-NN models by replacing the UQ methods for non-NN models such as [33, 34].
4 The Beauty Framework
To answer the questions in section 3, we propose our new framework for benefit estimation called Beauty (Benefit estimation with uncertainty).
4.1 Overview
The process of index tuning with Beauty is illustrated in Figure 6. After generating an index configuration by the enumerator, the advisor uses the uncertainty-aware learning-based BE model to estimate both the benefits that the configuration brings to the queries in the workload and the confidence levels of the estimated benefits. Then, the result filter examines the estimated benefits and their uncertainty with respect to a specified uncertainty threshold. The estimated benefits with uncertainty below the threshold are deemed reliable and are returned to the enumerator. For the remaining estimated benefits on which the BE model has low confidence due to insufficient training data or feature extraction deficiencies, the what-if tool that is usually more robust than learning-based BE models are invoked instead for benefit estimation.
Figure 6: The index tuning process with Beauty.
4.2 Uncertainty in Learning-based BE Models
The existing UQ methods are designed for general ML models and scenarios, lacking specific knowledge necessary for benefit estimation. To overcome this limitation, we propose a hybrid UQ method tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods. To this end, we conduct the first analysis of the sources of uncertainty in the BE problem and their relationships to the structures of learning-based BE models, which lays foundations for designing the uncertainty-aware BE model in Beauty.
Uncertainty emerges throughout the whole process of learning-based BE models from design to training and to inference. We divide the uncertainty into five categories which are summarized in Table I.
@@ 4.2 Uncertainty in Learning-based BE Models
U1: Epistemic uncertainty due to insufficient information modeling. Estimating the benefit of an index configuration for a query is a complex task, as it requires invoking the query optimizer to retrieve the plans of the query before and after creating the indexes in the configuration. Existing learning-based estimators only model the information most relevant to this query such as the index configuration, the current query plan, and the plan chosen by the optimizer after virtually creating the indexes in the configuration. However, some overlooked factors can also significantly affect query execution costs. One such factor is data distributions which determine the cardinalities of intermediate results produced during the execution of a query. Existing learning-based estimators either do not model data distributions or model them in an over-simplified manner. Cardinality estimation itself is challenging and is a key research focus in the AI4DB field. State-of-the-art data-driven cardinality estimation methods are often built on deep-learning models which are more complex than those used in the existing BE models. Another factor that significantly affects query execution costs is the configuration of the database system such as the knobs work_mem and shared_buffers. Finally, the implementation of the query optimizer itself can alter the selected plan for a query and varies its execution cost. Currently, none of the learning-based estimators has taken all these relevant information into account. The ignorance of these information leads to epistemic uncertainty.
@@
4.3 Hybrid Quantification Method
@@
4.3.1 Modules for Quantifying Uncertainty
As described in subsection 2.2, there are three main modules in a learning-based BE model: the feature extractor, the hidden vector encoder, and the predictor. Here, we analyze the relationship between the five categories of uncertainty listed in subsection 4.2 and the modules of the BE model, which lays the foundation for our hybrid UQ method.
@@ 4.3.1 Modules for Quantifying Uncertainty
4.3.2 Quantifying Uncertainty from Hidden Vector Encoder
@@
The rationale behind this approach lies in the fact that for the decoder to accurately reconstruct $V^{1}$ from $V^{h}$, two essential conditions must be satisfied. First, the structure of the hidden vector encoder must be appropriate. The network structure of the decoder should be symmetric to that of the encoder. Only if the encoder’s structure can gradually eliminate unimportant features and reduce dimensionality, the decoder can reverse this process and reconstruct $V^{1}$ based on the essential features. Second, the quality of the latent representation $V^{h}$ must be high. Given that the structure of the decoder is appropriate, the quality of the decoder’s input also determines the quality of the reconstruction. Therefore, when the divergence between $V^{1}$ and $\hat{V}^{1}$ is very small, the encoder has processed $V^{1}$ effectively, indicating low uncertainty in the hidden vector encoder module.
@@
4.3.3 Quantifying Uncertainty from Predictor
@@
[…]
The prior distribution $p(\mathbf{X}_{W})$ should reflect prior beliefs about the distribution of the weights $W$ before observing any data. Solving this equation directly can be challenging, so variational inference (VI) is often employed [36, 26]. Blundell et al. [31] proposed the renowned Bayes by Backprop (BBB) algorithm, which is compatible with backpropagation and facilitates learning a probability distribution. We select the BBB algorithm as the baseline in the subsequent experimental evaluation.
@@
(2) Monte Carlo Dropout (MCD) Methods. Dropout layers are commonly used in deep NNs to prevent overfitting. During training, a dropout layer randomly sets elements of the input tensor to 0 with probability $p\in[0,1]$ and scales up the remaining elements by $1/(1-p)$ to maintain the sum over all inputs. The zeroed elements are chosen independently for each forward pass. During inference, dropout layers are typically disabled. However, to represent model uncertainty, dropout layers can be activated during inference. Gal and Ghahramani [27] proved that dropouts can approximate Bayesian inference in BNNs. Given a test record $\mathbf{x}$, model inference is repeated multiple times independently, yielding outputs $\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}$. The mean of these outputs is the prediction for $\mathbf{x}$, and the variance represents the model’s uncertainty with respect to $\mathbf{x}$.
@@ 4.3.4 Result Filter
$\hat{B}(q,\mathcal{I}_{0},\mathcal{I})=\begin{cases}\hat{y}&\text{if }\mathcal{U}_{1}\leq\theta_{1}\ \text{and}\ \mathcal{U}_{2}\leq\theta_{2},\\ \text{What-if tools}&\text{otherwise}.\end{cases}$
The domain of $\theta_{1}$ and $\theta_{2}$ is $[0,+\infty)$, and their values vary based on factors such as model structure, hyperparameters, query workload, and the database instance. Setting fixed, universal thresholds for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ is impractical. Therefore, we adopt a context-dependent approach to define relative thresholds $\theta_{1}$ and $\theta_{2}$: for a new sample, if the model’s uncertainty regarding it is comparable to the uncertainty observed during training when the model performed optimally with minimal uncertainty, the result is deemed reliable. Specifically, after training the BE-UQ model, we perform inference for each training sample $(q,\mathcal{I}_{0},\mathcal{I})$ and gather the quantified uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$. The 90th percentile of $\mathcal{U}_{1}$ (or $\mathcal{U}_{2}$) across all training samples is designated as $\theta_{1}$ (or $\theta_{2}$).
4.4 Model Deployment
4.4.1 Model Training

### v2

# Can Uncertainty Quantification Improve Learned Index Benefit Estimation?

@@ I Introduction
(1) Poor stability: Learning-based models typically outperform what-if tools on data similar to their training sets, but they can generate highly erroneous results when query patterns shift or training data is insufficient. As shown in Figure 1, we compare a learning-based BE method with what-if tools on varying proportions of out-of-domain (OOD) samples. When the proportion reaches around 25%, the mean errors of two methods are comparable, but the mean error of the learning-based method becomes larger than the what-if tools as the OOD sample proportion increases.
@@
Due to the poor stability of learning-based models, some estimation results will inevitably be highly inaccurate and unreliable, leading the index advisor to select suboptimal indexes. Furthermore, the lack of interpretability prevents us from determining when or why the model produces unreliable results, making it difficult to fundamentally address the stability issue. In this paper, we propose a different approach by quantifying the uncertainty of the model’s results through specific metrics. When the model’s estimate for a test sample shows high uncertainty, the result is more likely to significantly deviate from the true value. This technique, known as uncertainty quantification (UQ), has been widely studied in fields where model errors have serious consequences, such as medical image detection, autonomous driving, and drug analysis [8, 9].
@@
(1) We conducted a systematic study of the sources of uncertainty in learning-based BE models within the context of index tuning. Based on the characteristics of these tasks, we proposed a novel hybrid UQ method that allows for a flexible trade-off between efficiency and accuracy, depending on the specific characteristics of the BE task (Section IV). Our experiments demonstrate that the proposed UQ method outperforms general-purpose UQ methods in five out of six scenarios.
(2) We introduced Beauty, the first uncertainty-aware BE framework, which combines the strengths of learning-based models with what-if tools, enabling uncertainty-driven model updates. Our experiments across six benchmarks with varying budgets show that the Beauty framework significantly improves index tuning performance (Section VI).
(3) In our experiments, we applied existing UQ methods to various learning-based BE models, creating a total of 16 different models (Section V). We thoroughly evaluated the impact of these methods on model accuracy and efficiency, confirming the superiority of our proposed UQ method.
II Background
II-A Index Advisors
@@
II-B Learning-based Benefit Estimation (BE) Models
The benefit estimator plays a crucial role in index tuning as its efficiency directly impacts the overall performance of index tuning, and its accuracy affects the behavior of the index configuration enumerator [14, 19, 20]. Traditionally, what-if tools have been used to create hypothetical indexes, and the query optimizer is employed to estimate the total cost $c(W,\mathcal{I})$. However, due to their inefficiency and inaccuracy, particularly, for complex queries, machine learning (ML) techniques have been recently introduced to develop faster and more accurate BE models. One approach such as DISTILL [1] involves training multiple lightweight BE models for each query template to collectively estimate benefits. The other approach is to train a single complex BE model using historical query execution data, either to completely replace the what-if tools or to adjust the outputs of the what-if tool. The latter approach is the main focus of our study. LIB [7] and AMA [6] are notable examples of this approach. As shown in Figure 3, the typical architecture of a learning-based BE model consists of three modules.
@@ II-B Learning-based Benefit Estimation (BE) Models
II-C Uncertainty in Machine Learning (ML) Models
Uncertainty is the quality or state of being uncertain, which arises in every stage of a ML task, including data acquisition, feature encoding, model design, model training, and model inference [8]. Naturally, uncertainty is inherent in ML-based index tuning, which will be explained in subsection IV-B. As illustrated in Figure 4, there are two different types of uncertainty, often referred to as aleatoric uncertainty and epistemic uncertainty [9].
@@ II-C Uncertainty in Machine Learning (ML) Models
Uncertainty Quantification (UQ): A number of model-independent and model-dependent methods have been proposed to quantify the two types of uncertainty. The first category includes Bayesian neural networks (BNNs) [22, 23], Monte Carlo dropout (MCD) methods [24], and ensemble methods [25], which do not rely on the specific models in which the uncertainty is quantified. The second category incorporates UQ into model-specific structures. For example, HSTO [26] replaces the multi-head attention mechanism with hierarchical stochastic self-attention to achieve UQ. In this paper, we will reproduce these methods in the experimental evaluation and compare them with our proposed hybrid UQ approach.
III Problem
Since collecting training data is costly, ML models inevitably encounter out-of-distribution (OOD) data, leading to unreliable outputs. The unreliable results produced by learning-based benefit estimators can lead index advisors to select inappropriate indexes, significantly degrading database performance. Therefore, as discussed in section I, we will explore how UQ techniques can be leveraged to determine whether the results produced by learning-based BE models are reliable and how to integrate uncertainty-aware benefit estimators into index advisors. Specifically, this paper addresses the following three key research questions:
Q1: How can the uncertainty in a learning-based BE model be precisely quantified? Accurately quantifying the uncertainty in a learning-based BE model presents a significant challenge, and directly applying the existing methods often yields suboptimal performance. As shown in Figure 5, on a OOD dataset, model error and model uncertainty are generally not highly linearly correlated, that is, a large model error may not mean large model uncertainty. In this paper, the first problem is how to precisely quantify the uncertainty in a learning-based BE model. Rather than applying the existing general-purpose UQ methods directly, we propose a hybrid quantification method in section IV tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods.
@@ III Problem
Q2: What impacts does UQ have on a BE model? In section V, we will conduct an empirical study to understand the impacts of UQ on BE models by evaluating the accuracy, training cost, and inference cost of the BE models enhanced with UQ. In subsection IV-D, we will also explore a new method for determining when a BE model needs to be updated based on quantified uncertainty.
Q3: What impacts does a UQ-enhanced BE model have on the index advisor? In section VI, we will evaluate 8 UQ-enhanced BE models with 6 workloads across 36 scenarios to investigate the effects of the accuracy and efficiency of UQ on the index advisor.
Due to space constraints, we focus exclusively on BE models based on neural networks (NNs). NN-based models have been receiving increasing attention and generally achieve higher accuracy. The approach proposed in this paper can be easily extended to non-NN models by replacing the UQ methods for non-NN models such as [28].
IV The Beauty Framework
To answer the questions in section III, we propose our new framework for benefit estimation called Beauty (Benefit Estimation with UncertainTY).
IV-A Overview
The index tuning process with Beauty is illustrated in Figure 6. After the enumerator generates an index configuration, the advisor uses the uncertainty-aware learning-based BE model to estimate both the benefits that the index configuration brings to the queries in the workload and the uncertainties of the estimated benefits. Then, the result filter examines the estimated benefits and their uncertainties with respect to a specified uncertainty threshold. The estimated benefits with uncertainties below the threshold are deemed reliable and returned to the enumerator. For the remaining estimated benefits with uncertainties above the threshold due to insufficient training data or deficiencies in feature extraction, the what-if tool is invoked for benefit estimation because it is usually more robust than learning-based BE models.
Figure 6: Index tuning process with Beauty.
IV-B Uncertainties in Learning-based BE Models
The existing UQ methods are designed for general ML models and scenarios, lacking specific knowledge necessary for BE. To overcome this limitation, we propose a hybrid UQ method tailored to the unique characteristics of the BE problem, providing more precise quantification of uncertainty than the general-purpose methods. To this end, we conducted the first analysis of the sources of uncertainties in the BE problem and their relationships with the structures of learning-based BE models, which lays the foundations for designing the uncertainty-aware BE model in Beauty.
Uncertainties emerge throughout the entire process of learning-based BE models, from model design to model training, and to model inference. Uncertainties can be divided into five categories as summarized in Table I.
@@ IV-B Uncertainties in Learning-based BE Models
U1: Epistemic uncertainty due to insufficient information modeling. Estimating the benefit of an index configuration for a query is a complex task, as it requires invoking the query optimizer to retrieve the plans of the query before and after creating the indexes in the configuration. Existing learning-based estimators only model the information most relevant to this query, such as the index configuration, the current query plan, and the plan chosen by the optimizer after virtually creating the indexes in the configuration. However, some overlooked factors can also significantly affect query execution costs. One such factor is data distributions that determine the cardinalities of intermediate results produced during the execution of a query. Existing learning-based estimators either do not model data distributions or model them in an over-simplified manner. Cardinality estimation itself is challenging and is a key research focus in the AI4DB field. State-of-the-art data-driven cardinality estimation methods are often built on deep learning models that are more complex than those used in existing BE models. Another factor that significantly affects query execution costs is the configuration of the database system such as the knobs work_mem and shared_buffers. Finally, the implementation of the query optimizer itself can alter the selected plan for a query and vary its execution cost. Currently, none of the learning-based estimators has taken all this relevant information into account. The ignorance of related information leads to epistemic uncertainty.
@@
IV-C Hybrid Quantification Method
@@
IV-C1 Modules for Quantifying Uncertainty
As described in subsection II-B, there are three main modules in a learning-based BE model: the feature extractor, the hidden vector encoder, and the predictor. Here, we analyze the relationship between the five categories of uncertainty listed in subsection IV-B and the modules of the BE model, which lays the foundation for our hybrid UQ method.
@@ IV-C1 Modules for Quantifying Uncertainty
IV-C2 Quantifying Uncertainty from Hidden Vector Encoder
@@
The rationale behind this approach lies in the fact that for the decoder to accurately reconstruct $V^{1}$ from $V^{h}$, two essential conditions must be satisfied. First, the mechanism and architecture implemented by the hidden vector encoder must be well-designed. The decoder typically employs identical mechanisms with a mirrored network structure. Only if the encoder’s structure can gradually eliminate unimportant features and reduce dimensionality, the decoder can reverse this process and reconstruct $V^{1}$ based on the essential features. Second, the quality of the latent representation $V^{h}$ must be high. Given that the structure of the decoder is appropriate, the quality of the decoder’s input also determines the quality of the reconstruction. Therefore, when the divergence between $V^{1}$ and $\hat{V}^{1}$ is very small, the encoder has processed $V^{1}$ effectively, indicating low uncertainty in the hidden vector encoder module.
@@
IV-C3 Quantifying Uncertainty from Predictor
@@
$=\frac{p(D|\mathbf{X}_{W})p(\mathbf{X}_{W})}{\idotsint p(D|\mathbf{X}_{W})p(\mathbf{X}_{W})d\mathbf{X}_{W}}.$
The prior distribution $p(\mathbf{X}_{W})$ should reflect prior beliefs about the distribution of the weights $W$ before observing any data. Solving this equation directly can be challenging, so variational inference (VI) is often employed [23]. Blundell et al. [27] proposed the renowned Bayes by Backprop (BBB) algorithm, which is compatible with backpropagation and facilitates learning a probability distribution. We select the BBB algorithm as the baseline in the subsequent experimental evaluation.
@@
(2) Monte Carlo Dropout (MCD) Methods. Dropout layers are commonly used in deep NNs to prevent overfitting. During training, a dropout layer randomly sets elements of the input tensor to 0 with probability $p\in[0,1]$ and scales up the remaining elements by 1/(1p)1/(1-p) to maintain the sum over all inputs. The zeroed elements are chosen independently for each forward pass. During inference, dropout layers are typically disabled. However, to represent model uncertainty, dropout layers can be activated during inference. Gal and Ghahramani [24] proved that dropouts can approximate Bayesian inference in BNNs. Given a test record $\mathbf{x}$, model inference is repeated multiple times independently, yielding outputs $\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{m}$. The mean of these outputs is the prediction for $\mathbf{x}$, and the variance represents the model’s uncertainty with respect to $\mathbf{x}$.
@@ IV-C4 Result Filter
$\hat{B}(q,\mathcal{I}_{0},\mathcal{I})=\begin{cases}\hat{y_{1}}&\text{if }\mathcal{U}_{1}\leq\theta_{1}\ \text{and}\ \mathcal{U}_{2}\leq\theta_{2},\\ \hat{y_{2}}&\text{otherwise},\end{cases}$
where $\hat{y_{1}}$ and $\hat{y_{2}}$ represent the results obtained from the learning-based model and the what-if tool, respectively. The domain of $\theta_{1}$ and $\theta_{2}$ is $0,+\infty)$, and their values are influenced by factors such as model architecture, hyperparameters, query workload, and the database instance. Consequently, it is impractical to set fixed, universal thresholds for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$. However, we observe two key facts: (1) A well-trained learning-based model yields accurate and reliable results on the training dataset, where uncertainty is minimized. (2) The uncertainties quantified for the training instances typically follow a long-tailed distribution, with over 94% of cases in the 54 experimental results presented in subsection V-C conforming to this trend.
Based on these observations, we propose a context-aware method to determine $\theta_{1}$ and $\theta_{2}$. After training the BE-UQ model, we infer the BE-UQ model for each training sample $(q,\mathcal{I}_{0},\mathcal{I})$ and collect the uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ quantified for $(q,\mathcal{I}_{0},\mathcal{I})$. A naïve method is to set $\theta_{i}=\max(\mathcal{U}_{i})$. However, due to the long-tailed distribution of $\mathcal{U}_{i}$, the presence of a few large uncertainties (outliers) can make $\theta_{i}$ unnecessarily large, causing many unreliable samples not identified. Therefore, we refer to the Interquartile Range (IQR) method [30] to set
$\theta_{i}=\min(P_{75}(\mathcal{U}_{i})+\alpha\times(P_{75}(\mathcal{U}_{i})-P_{25}(\mathcal{U}_{i})),\max(\mathcal{U}_{i})),$
where $P_{25}(\mathcal{U}_{i})$ and $P_{75}(\mathcal{U}_{i})$ are the 25th and 75th percentiles of the collected values of $\mathcal{U}_{i}$, respectively. The parameter $\alpha$ is typically set within the range of $1$ to $1.5$. Due to the long-tail distribution of $\mathcal{U}_{i}$, we usually have $\theta_{i}\ll\max(\mathcal{U}_{i})$, which can significantly increase the recall of unreliable samples.
IV-D Model Deployment
IV-D1 Model Training